Please show all work.

2. Two parallel transmission lines are 75cm apart and carry currents of 11 35A and I2 -15A. (a) (5 points) Draw the magnetic field around each wire. I2 (b) (10 points) At what distance, r, from the top wire is the net magnetic field of the two wires zero 1 35 A d75 cm 12 15 A
Answer #1

a) magnetic field lines according to Fleming right hand rule

in between wires magnetic field lines are opposite in direction and circular around wires,

means attractive in nature.

b) let at a distance r from wire 1 magnetic field will be zero.

Bnet = Uo*I1/(2*pi*r) - Uo*I2/(2*pi*(d-r)) = 0

=> I1/r - I2/(d-r) = 0

=> (d-r)/r = I2/I1 = 15/35 = 3/7

=> 7d = 10r

=> r = 7*75/10 = 52.5 cm